OpenStudy (anonymous):

Solve for x: 4 over 5 x + 4 over 3 = 2x x = __________________ Write your answer as a fraction in simplest form. Use the "/" symbol for the fraction bar.

OpenStudy (anonymous):

Medal who gives the correct answer! :D :D :D :D

OpenStudy (anonymous):

10 over 9 I believe

OpenStudy (anonymous):

10/9

OpenStudy (anonymous):

(4/5)*x+4/3 = 2*x // - 2*x (4/5)*x-(2*x)+4/3 = 0 (4/5)*x-2*x+4/3 = 0 4/3-6/5*x = 0 // - 4/3 -6/5*x = -4/3 // : -6/5 x = -4/3/(-6/5) x = 10/9 x = 10/9

OpenStudy (ahsome):

\[x=\frac{10}{9}\] According to @caleb26147 :)

OpenStudy (ahsome):

\[\frac{4}{5x}+\frac{4}{3}=2x\] Is this the question? @horselover33